# Some characterizations of weakly uniformly perfect sets

### v5

@@ 1. Introduction
$(1)$ ()($\Rightarrow$)( ⇒ ): Write $h(r)=C_{0}h_{1,\alpha}(r)=C_{0}r^{\alpha}$. Then there exists a constant $r_{0}>0$, such that for every $a\in E$ and $r\in(0,r_{0})$, we can construct a Cantor-type set $\mathcal{C}\subset E\cap\overline{B}(a,r)$ as above. Let $A=\overline{B}(x,\rho)$.
@@
Notice that these discs do not inersect each other. Therefore, we see that
#$\#J\cdot\pi\cdot(\tilde{h}^{\circ k}(r))^{2}\leq\pi(\rho+2\tilde{h}^{\circ k})^{2}.$
@@
#$\#J\leq\left(\frac{\rho}{\tilde{h}^{\circ k}(r)}+2\right)^{2}\leq 9.$
Thus we have
@@
where $C_{2}=C_{1}^{\frac{1}{\alpha-1}}$. Thus for
@@
Now for any countable cover $\{A_{j}\}$ of $\mathcal{C}$ by closed balls $A_{j}$ are discs with radii $\rho_{j}$, we have
@@
Since $\{A_{j}\}$ is arbitrary, we have
@@
Since $a\in E,r\in(0,r_{0})$ are arbitrary, we obtain the necessity of $(1)$.
()($\Leftarrow$)( ⇐ ): Now for every $a\in E$ and $r\in(0,r_{0})$, we have
@@
Let $h(t)=t^{\alpha}$, here $\alpha>1$ is undetermined. Then
@@
We claim that there are $\alpha>1$ and $0<r_{1}\ll 1$, such that
@@
To illustrate this statement, we only need to show
@@
By claim, (5.4) and (5.5), for every $a\in E$ and $r\in(0,r_{1})$, we have
@@
$(2)$ ()($\Rightarrow$)( ⇒ ) The proof is basically the same as $(1)$, except that the form of (5.3) may change due to the change in the selection of $h$. Now we have $\tilde{h}(r)=C_{0}r\left(\log\frac{2}{r}\right)^{-\beta}$ for some $C_{0}>0$. Write $s_{0}=r,s_{k+1}=\tilde{h}(s_{k})$. In view of the derivation process of (3.25) in [34], there exist a $C_{1}>0$ such that
@@
on $(0,2r_{0})$. Replace (5.3) in $(1)$ with (5.8) to obtain the necessity part of $(2)$.
()($\Leftarrow$)( ⇐ ) Similar to $(1)$, we have
@@
We claim that there exist $\beta>0$ and $0<r_{2}\ll 1$, such that
@@
In fact, we only need to show
@@
So the claim is true if we choose $\beta\gg 1$ and $r_{2}\ll 1$. The proof is concluded as in $(1)$. ∎
@@
There exists a compact set $E\subset\mathbb{C}$, such that $\dim_{H}(E)=0$ and $E$ satisfies condition $(U)_{1,\alpha}$ (or condition $(U)_{2,\beta}$).
Consider Cantor-type set $\mathcal{C}$: choose a sequence $\{l_{j}\}_{j=0}^{\infty}$, where $l_{j}>0,l_{j+1}<\frac{l_{j}}{2}$. Set $\mathcal{C}_{0}=[0,l_{0}]$ and define $\mathcal{C}_{j}$ to be a union of $2^{j}$ closed intervals inductively, such that $\mathcal{C_{j}}$ is obtained by removing from the middle of each inteval in $\mathcal{C}_{j-1}$ an open subinteval whose length is $l_{j-1}-2l_{j}$. For example, $\mathcal{C}_{1}=[0,l_{1}]\cup[l_{0}-l_{1},l_{0}]$, $\mathcal{C}_{2}=[0,l_{2}]\cup[l_{1}-l_{2},l_{1}]\cup[l_{0}-l_{1},l_{0}-l_{1}+l_{2}]\cup[l_{0}-l_{2},l_{0}]$, etc. Write
@@
where every $I_{j,k}$ is a closed inteval of length $l_{j}$, lying on the left of $I_{j,k+1}$. We set
@@
$\Gamma^{\gamma}(\mathcal{C})\leq\Gamma^{\gamma}(\mathcal{C}_{j})\leq\sum_{k=1}^{2^{j}}l_{j}^{\gamma}=2^{j}\cdot l_{j}^{\gamma}.$
@@
$00\leq\Gamma^{\gamma}(\mathcal{C})\leq\lim_{j\to\infty}\Gamma^{\gamma}(\mathcal{C}_{j})=$ 0
@@
Similarly, if we take $l_{j}=l_{j-1}\left(\log\frac{1}{l_{j}}\right)^{-\beta}<\frac{l_{j}}{2}$, then the corresponding $\mathcal{C}$ satisfies $(U)_{2,\beta}$ condition and $dim_{H}(\mathcal{C})=0$.

### v6

@@ 1. Introduction
$(1)$ $(\Rightarrow)$: Write $h(r)=C_{0}h_{1,\alpha}(r)=C_{0}r^{\alpha}$. Then there exists a constant $r_{0}>0$ such that for every $a\in E$ and $r\in(0,r_{0})$, we can construct a Cantor-type set $\mathcal{C}\subset E\cap\overline{B}(a,r)$ as above. Let $A=\overline{B}(x,\rho)$.
@@
Notice that these discs do not inersect each other. Therefore, we observe that
$J\cdot\pi\cdot(\tilde{h}^{\circ k}(r))^{2}\leq\pi(\rho+2\tilde{h}^{\circ k})^{2}.$
@@
$J\leq\left(\frac{\rho}{\tilde{h}^{\circ k}(r)}+2\right)^{2}\leq 9.$
Thus, we have
@@
where $C_{2}=C_{1}^{\frac{1}{\alpha-1}}$. Thus, for
@@
Now for any countable cover $\{A_{j}\}$ of $\mathcal{C}$ by closed discs $A_{j}$ with radii $\rho_{j}$, we have
@@
Since $\{A_{j}\}$ is arbitrary, we obtain
@@
Since $a\in E,r\in(0,r_{0})$ are arbitrary, we conclude the necessity of $(1)$.
$(\Leftarrow)$: Now, for every $a\in E$ and $r\in(0,r_{0})$, we have
@@
Let $h(t)=t^{\alpha}$, where $\alpha>1$ is undetermined. Then
@@
We claim that there exist $\alpha>1$ and $0<r_{1}\ll 1$ such that
@@
To establish this claim, it suffices to show that
@@
By the claim, inequalities (5.4) and (5.5), for every $a\in E$ and $r\in(0,r_{1})$, we have
@@
$(2)$ $(\Rightarrow)$ The proof is essentially the same as in $(1)$, except that the form of (5.3) may change due to the change in the choice of the function $h$. Now, we have $\tilde{h}(r)=C_{0}r\left(\log\frac{2}{r}\right)^{-\beta}$ for some constant $C_{0}>0$. Write $s_{0}=r,s_{k+1}=\tilde{h}(s_{k})$. In view of the derivation process of (3.25) in [34], there exists a constant $C_{1}>0$ such that
@@
on $(0,2r_{0})$. Replacing (5.3) in $(1)$ with (5.8) to obtain the necessity part of $(2)$.
$(\Leftarrow)$ Similar to $(1)$, we have
@@
We claim that there exist constants $\beta>0$ and $0<r_{2}\ll 1$, such that
@@
In fact, it suffices to show that
@@
So the claim is verified if we choose $\beta\gg 1$ and $r_{2}\ll 1$. The proof is concluded as in $(1)$. ∎
@@
There exists a compact set $E\subset\mathbb{C}$ such that $\dim_{H}(E)=0$ and $E$ satisfies condition $(U)_{1,\alpha}$ (or condition $(U)_{2,\beta}$).
Consider a Cantor-type set $\mathcal{C}$: choose a sequence $\{l_{j}\}_{j=0}^{\infty}$, where $l_{j}>0$ and $l_{j+1}<\frac{l_{j}}{2}$. Set $\mathcal{C}_{0}=[0,l_{0}]$ and define $\mathcal{C}_{j}$ to be a union of $2^{j}$ closed intervals inductively, such that $\mathcal{C_{j}}$ is obtained by removing from the middle of each inteval in $\mathcal{C}_{j-1}$ an open subinteval of length $l_{j-1}-2l_{j}$. For example, $\mathcal{C}_{1}=[0,l_{1}]\cup[l_{0}-l_{1},l_{0}]$, $\mathcal{C}_{2}=[0,l_{2}]\cup[l_{1}-l_{2},l_{1}]\cup[l_{0}-l_{1},l_{0}-l_{1}+l_{2}]\cup[l_{0}-l_{2},l_{0}]$, etc. Write
@@
where each $I_{j,k}$ is a closed inteval of length $l_{j}$, lying on the left of $I_{j,k+1}$. We set
@@
$\Lambda^{\gamma}(\mathcal{C})\leq\Lambda^{\gamma}(\mathcal{C}_{j})\leq\sum_{k=1}^{2^{j}}l_{j}^{\gamma}=2^{j}\cdot l_{j}^{\gamma}.$
@@
$0\leq\Lambda^{\gamma}(\mathcal{C})\leq\lim_{j\to\infty}\Lambda^{\gamma}(\mathcal{C}_{j})=0$
@@
Similarly, if we take $l_{j}=l_{j-1}\left(\log\frac{1}{l_{j}}\right)^{-\beta}<\frac{l_{j}}{2}$, then the corresponding $\mathcal{C}$ satisfies condition $(U)_{2,\beta}$ and $dim_{H}(\mathcal{C})=0$.